# CodePlan: Unlocking Reasoning Potential in Large Language Models by Scaling Code-form Planning

### v1

@@ 3.4 Evaluation Benchmarks
Mathematical Reasoning. We employ three benchmark datasets of mathematical problems: (1) GSM8K (Cobbe et al., 2021), a collection of grade-school problems; (2) MATH (Hendrycks et al., 2021), a more challenging suite of high-school-level problems; and (3) SVAMP (Patel et al., 2021), a robustness evaluation benchmark.
Symbolic Reasoning. We use four benchmarks that require multi-step logical deductions: (1) Boolean Expression from Big-bench-hard (BBH) (Suzgun et al., 2022), where the model infers the final value of a complex boolean expression; (2) Coin Flipping (Wei et al., 2022b), which requires determining the ultimate face after a sequence of coin flips. We use a more challenging 4-flip version during evaluation; (3) Last Letter Concatenation (Wei et al., 2022b), which requires concatenating the final letters of a word sequence, also evaluated with a more challenging 4-word version; and (4) Dyck Language from BBH, which requires predicting the sequence of the closing parentheses of a Dyck-4 word without its last few closing parentheses. The Dyck Language task often demands over 10 reasoning steps, making it particularly challenging. Under few-shot evaluations, we observe that all models struggle to achieve non-trivial accuracy due to the degeneration phenomena (Holtzman et al., 2020), an inherent limitation of LLMs in long text generation. To mitigate these confounding factors beyond planning capabilities, we continue fine-tuning each trained model on 5K examples that are automatically constructed using official implementation before evaluation.
@@ 3.5 Main Results
As highlighted in Table 3, CodePlan demonstrates consistent and substantial performance improvements across all base models and most benchmarks against the baselines, underscoring the efficacy of incorporating code-form plans in enhancing LLMs’ systematic reasoning capabilities. Specifically, CodePlan generally outperforms the PS Prompting baseline, often by a substantial margin. This indicates that the benefits of learning to plan in code are beyond what can be achieved by planning in natural language with careful prompt engineering alone. Notably, the vanilla training approach, which solely learns low-level reasoning steps, does not always improve backbone models’ performance on downstream reasoning tasks. For instance, the performance of Mistral-7B drops by $2$ to $9$ points on several symbolic, multi-hop QA and decision-making tasks after vanilla training. We attribute this to the distribution gap between the massive, wide-ranging corpus and downstream benchmarks, thereby leading to suboptimal adaptation of the model’s reasoning capabilities. In contrast, CodePlan consistently improves upon the initial backbone model across all benchmarks, demonstrating its ability to develop more generalizable and transferable planning knowledge during training. The superiority may result from the inherent structured, concise, and unambiguous semantics encoded within code-form plans. In this way, LLMs can more effectively extract and internalize the underlying planning signals in natural language data.
@@ 3.6 Analysis
In this section, we present an in-depth analysis regarding two key benefits of CodePlan: (1) increasing superiority on more complex problems (Finding 1) and (2) improved training data efficiency (Finding 2). Furthermore, we compare CodePlan against two variants, including planning in natural language (Finding 3) and directly reasoning through executable code (Finding 4), offering empirical evidence for the merits of leveraging codes to represent intermediate plans. Appendix B also discusses the influence of plan annotation models and the efficiency of CodePlan, and presents several cases for illustrating the superiority and limitations of CodePlan.
@@
To gain deeper insights into the merits of CodePlan, we conduct experiments on the MusiQue dataset, which encompasses questions spanning various levels of complexity. As illustrated in Figure 2, we observe that the relative performance gains yielded by CodePlan become more pronounced as the complexity of the tasks increases. The relative improvements in EM scores increase from 6.3% for 2-hop questions to a remarkable 43.8% for the most challenging 4-hop questions.
This finding highlights a key insight – as reasoning challenges grow more intricate and multi-faceted, the ability to generate and leverage structured code-form plans becomes increasingly beneficial. In simple tasks, LLMs’ inherent language understanding capabilities are often sufficient. However, as task complexity increases, the limitations of vanilla training—the ambiguity and obfuscation of planning signals in natural language data—become more apparent. In such scenarios, CodePlan enables LLMs to systematically understand and frame the reasoning process, thereby navigating solution pathways more effectively.
Figure 3: Performance trajectories on two downstream tasks during vanilla training and CodePlan. “Musique 4-Hop” means the models are evaluated on only 4-hop questions of the Musique benchmark.
@@
We analyze the performance trajectories of LLMs during training with or without CodePlan. As depicted in Figure 3, the LLM trained with CodePlan almost always outperforms its counterpart trained on the same raw prompt-response data, evincing superior knowledge acquisition and generalization to out-of-distribution reasoning challenges. Moreover, CodePlan achieves a more stable and consistent performance ascent, maintaining its supremacy throughout the training process. This showcases the advantage of CodePlan in developing transferable high-level reasoning skills from wide-ranging data, as illustrated by the case in Appendix B.3.
Figure 4: Comparing planning in natural language with CodePlan. The scores of each type of task are calculated by averaging the metric scores across all corresponding benchmarks, except multi-hop QA tasks where we average the EM scores.
@@
To substantiate the merits of encoding planning knowledge through code versus natural language, we conduct comparative experiments by replacing the code plans in our dataset with natural language counterparts. Specifically, we generate natural language plans following the same pipeline in §3.1, while using a modified prompt shown in Appendix A.3.
Figure 4 shows the results. Training with natural language plans also leads to moderate improvement over the vanilla baseline. However, CodePlan consistently outperforms its natural language counterparts across all benchmarks by substantial margins. The performance gaps are particularly pronounced on complex reasoning tasks requiring structured planning, such as mathematical reasoning, symbolic reasoning, instruction-following, and decision-making tasks, with relative improvements of 4%, 27.2%, 6.5%, and 27.5%, respectively. These findings underscore the superiority of code representations over natural language counterparts.
Method Stage 1 Stage 2 Overall
$-\log p(Z^{*}|X)$ $-\log p(Y|X,Z^{*})$
@@
Table 4: Negative log-likelihood of generating responses using different plan forms. Stage 1 and Stage 2 mean the planning stage and surface realization stage, respectively. “Overall” is calculated by summing up the two parts.
To gain further insights into how planning in code outperforms natural language, we evaluate the performance of the two-stage process: planning and surface realization. For each plan form, Table 4 reports the model’s negative log-likelihood (NLL) on a subset of the training data (10K examples). The vanilla training baseline, which forgoes explicit planning, exhibits a high overall NLL score, underscoring its inherent struggle to directly model the complex mapping from prompts to final responses. While incorporating natural language plans substantially reduces the NLL for the surface realization stage, it introduces a significant challenge in planning compared to CodePlan. We attribute this to two reasons: (1) code provides a more structured and precise representation of complex reasoning logic compared to natural language, thus offering more concise and easier-to-learn plan labels. (2) by framing plan generation into code generation, LLMs can benefit more from the abilities absorbed during training, as code is much more prevalent than natural language plans in the pre-training corpora. Consequently, CodePlan achieves the most substantial overall NLL improvement, providing empirical validation for our theoretical framework in Eq. 4 that requires minimizing the combined difficulty of planning and surface realization.
@@
Previous work has primarily focused on leveraging executable code to directly solve mathematical reasoning tasks (Gao et al., 2023). However, we posit that this approach is inherently limited when tackling broad, multi-step reasoning challenges that necessitate deep natural language understanding capabilities. To substantiate this claim, we conduct comparative experiments by replacing the code plans in our dataset with executable code solutions. Specifically, instead of constructing pseudo-code plans, we directly translate each response into executable code that can output the answer after execution with a code interpreter, using the prompt in Appendix A.3. We refer to this baseline as CodeReason.
As presented in Figure 5, while CodeReason achieves the highest performance on SVAMP, a benchmark comprised of relatively simple math word problems, it consistently underperforms CodePlan across a diverse array of reasoning benchmarks that demand extensive natural language understanding capabilities, such as MATH, MuSiQue, and HotpotQA. Notably, CodeReason even lags behind the vanilla training baseline on MATH and MuSiQue. In contrast, CodePlan’s novel paradigm of generating code plans as an intermediate representation seamlessly integrates robust planning capabilities with the rich language understanding abilities of LLMs, yielding superior performance on intricate multi-step reasoning challenges.
@@ 4 Related Work
Endowing LLMs with robust reasoning abilities remains a formidable challenge. Existing approaches predominantly fall into three categories: (1) Prompting techniques leverage expert-designed prompts and demonstrations to elicit LLMs’ reasoning skills without additional training (Wei et al., 2022b; Press et al., 2023; Imani et al., 2023; Hong et al., 2024) (2) Task-specific fine-tuning aims to enhance LLMs’ reasoning capabilities on specific tasks by curating tailored fine-tuning data or rewards, spanning domains such as mathematical reasoning (Yu et al., 2023; Mitra et al., 2024; Shao et al., 2024; OpenAI, 2024), code reasoning (Le et al., 2022; Shen et al., 2023), instruction-following (Cui et al., 2023), visual reasoning (Cheng et al., 2024), and decision-making (Zeng et al., 2023; Guan et al., 2024) tasks. However, these task-specific approaches often falter in generalizing beyond their intended scopes. Very recently, OpenAI’s o1-preview (OpenAI, 2024) has exhibited remarkable performance in mathematical and code reasoning tasks. Nonetheless, it has also encountered skepticism regarding its extremely low inference efficiency and limited generalizability to other domains. Besides, the multi-path reasoning and reflection capabilities demonstrated by this model are orthogonal to our work, which we will explore in future research. (3) Tool integration seeks to augment LLMs’ reasoning capabilities by seamlessly incorporating external tools such as calculators (Schick et al., 2024), retrievers (Asai et al., 2023), and code interpreters (Gao et al., 2023) into the reasoning process, despite the inherent shortcomings of LLMs in utilizing diverse tools in concert (Wang et al., 2024).
@@
Enhancing LLMs’ planning capabilities is crucial for solving complex reasoning tasks (Yang et al., 2023). Existing approaches broadly fall into three groups: (1) Prompting for planning, which elicits planning skills through advanced prompting during inference (Wei et al., 2022a; Wang et al., 2023a; Khot et al., 2022); (2) Expert-guided planning, which trains LLMs based on human-curated, task-specific plans (Yin et al., 2024; Guan et al., 2024); and (3) Implicit planning, which models planning as latent variables to capture the high-level aspects when generating the response (Kingma, 2013). Although the implicit encoding approach alleviates the limitations of explicit plans in scalability, generalization, and reliance on expert knowledge, it may suffer from two issues: (a) the expressiveness of implicit plans may be limited by the feasible space of the chosen posterior distribution, often leading to the issue of posterior collapse (Bowman et al., 2016); and (b) it necessitates online sampling from both prior and posterior distributions during training, which can substantially impact computational efficiency. In contrast, we propose a scalable paradigm of explicitly teaching LLMs to plan for reasoning from large pre-training corpora, which is distinguished by leveraging an expressive and interpretable code representation for planning, rigorously demonstrating our approach’s efficacy.
@@
Recent works have explored leveraging code to empower LLMs for complex reasoning. The most direct approach is to employ executable code as a replacement for natural language in representing reasoning steps (Gao et al., 2023; Ye et al., 2023), albeit facing challenges when extending to broader, more complex reasoning domains. Beyond direct problem-solving, code has also been utilized to enhance LLMs’ capabilities in handling structured reasoning tasks, such as graph generation (Madaan et al., 2022), event structure prediction (Wang et al., 2022), and decision-making (Wang et al., 2024) tasks. Sharing a similar motivation, our work capitalizes on code’s superiority in representing intricate reasoning structures, introducing it as an intermediate plan to guide low-level reasoning steps across a wide range of tasks.
@@ 5 Conclusion
In this work, we introduce a pioneering paradigm to endow LLMs with robust planning capabilities through the explicit supervision of code-form plans. By reframing plan generation as a code generation task, our approach, termed CodePlan, harnesses the structured and versatile nature of programming languages to capture the rich semantics and control flows underpinning sophisticated reasoning processes. Importantly, CodePlan allows the automatic extraction of code-form plans from massive, wide-ranging text corpora without the need for curated, task-specific datasets. We demonstrate the effectiveness of this paradigm by training CodePlan on a large-scale dataset comprised of 2M natural language problems paired with their corresponding code-form plans. Across an extensive evaluation spanning 13 challenging multi-step reasoning benchmarks, CodePlan demonstrates remarkable efficacy, consistently and substantially outperforming vanilla training by a substantial margin. In-depth analyses further corroborate CodePlan’s increasing performance gain on complex problems and generalization ability. This work paves the way for several promising research directions, including exploring diverse posterior distributions over plans, empowering LLMs to generate code to drive low-level reasoning, and realizing agents that can seamlessly leverage external knowledge sources and APIs within their planning and reasoning processes.
@@ A.1 Training Data Curation Details
The curation of high-quality training data is crucial for the success of our approach. To ensure the integrity and relevance of the generated code-form plans, we filter examples without valid terminations in the plans, as this is essential for maintaining logical coherence.
@@ A.2 Data Statistics
Table 5: Statistics of the curated training data.
@@
Table 6: Number of test examples in the evaluation benchmarks.

### v2

@@ 3.4 Evaluation Benchmarks
Mathematical Reasoning. It involves three benchmarks: (1) GSM8K (Cobbe et al., 2021), a collection of grade-school problems; (2) MATH (Hendrycks et al., 2021), a more challenging suite of high-school-level problems; and (3) SVAMP (Patel et al., 2021), a robustness evaluation benchmark.
Symbolic Reasoning. We use four benchmarks requiring multi-step logical deductions: (1) Boolean Expression from Big-bench-hard (BBH) (Suzgun et al., 2022), where the model infers the value of a complex boolean expression; (2) Coin Flipping (Wei et al., 2022b), which requires determining the final face after a sequence of flips. We use the challenging 4-flip version; (3) Last Letter Concatenation (Wei et al., 2022b), which requires concatenating the last letter of a word sequence, evaluated with a 4-word version; and (4) Dyck Language from BBH, which requires predicting the closing parentheses of a Dyck-4 word. This task often demands over 10 reasoning steps. Under few-shot evaluations, we observe that all models struggle to achieve non-trivial accuracy due to the degeneration phenomena (Holtzman et al., 2020), an inherent limitation of LLMs in long text generation. To mitigate these confounding factors beyond planning capabilities, we synthesize 5K examples based on official implementation to fine-tune each model before evaluation.
@@ 3.5 Main Results
As highlighted in Table 3, CodePlan shows consistent and substantial performance improvements across all backbone models and most benchmarks against the baselines, underscoring the efficacy of incorporating code-form plans in enhancing LLMs’ systematic reasoning capabilities. Specifically, CodePlan generally outperforms the PS Prompting baseline, often by a substantial margin. This indicates that the benefits of learning to plan in code are beyond what can be achieved by planning in natural language with careful prompt engineering alone. Quiet-STaR demonstrates inferior performance even compared to the backbone model despite significant computation overhead [^2] We use the official implementation for Quiet-STaR, which only supports Mistral and not Llama series., illustrating the difficulty of learning implicit plans through latent variables. Notably, the vanilla training approach, which solely learns low-level reasoning steps, does not always improve backbone models’ performance on downstream tasks. For instance, the performance of Mistral-7B drops by $2$ to $9$ points on several symbolic, multi-hop QA and decision-making tasks after vanilla training. We attribute this to the distribution gap between the massive, wide-ranging corpus and downstream benchmarks, thereby leading to suboptimal adaptation of the model’s reasoning capabilities. In contrast, CodePlan consistently improves upon the initial backbone model across all benchmarks, demonstrating its ability to develop more generalizable and transferable planning knowledge during training. The superiority may result from the inherent structured, concise, and unambiguous semantics encoded within code-form plans. In this way, LLMs can more effectively extract and internalize the underlying planning signals in natural language data.
@@ 3.6 Analysis
We analyze two key benefits of CodePlan: (1) increasing superiority on more complex problems (Finding 1) and (2) improved training data efficiency (Finding 2). Furthermore, we compare CodePlan against two variants: planning in natural language (Finding 3) and reasoning through executable code (Finding 4), offering empirical evidence for the merits of leveraging code as intermediate plan representations. Appendix B also discusses the influence of plan annotation models, the efficiency of CodePlan, and case studies highlighting the strength and limitations of CodePlan.
@@
To gain deeper insights into the merits of CodePlan, we conduct experiments on MusiQue, which encompasses questions spanning various levels of complexity. As illustrated in Figure 2, CodePlan yields increasing performance gain as the task becomes more complex. The relative improvements in EM scores increase from 6.3% for 2-hop questions to a remarkable 43.8% for 4-hop questions.
This finding highlights a key insight – as reasoning challenges grow more intricate, the ability to generate and leverage structured code-form plans grows more valuable. For simple tasks, LLMs’ inherent language understanding capabilities often suffice. But as task complexity increases, the limitation of vanilla training—the ambiguity and obfuscation of planning signals in natural language data—becomes more clear. In such scenarios, CodePlan enables LLMs to systematically understand and frame the reasoning process, thereby navigating solution pathways more effectively.
Figure 3: Performance trajectories on two downstream tasks via vanilla training and CodePlan. “4-Hop” denotes evaluating on the 4-hop subset.
@@
We analyze the performance trajectories of CodePlan and vanilla training. As depicted in Figure 3, the LLM trained with CodePlan almost always outperforms its counterpart trained on the same raw prompt-response data, evincing superior knowledge acquisition and generalization to out-of-distribution reasoning challenges. Moreover, CodePlan achieves a more stable and consistent performance ascent, maintaining its supremacy throughout the training process. This showcases the advantage of CodePlan in developing transferable high-level reasoning skills from wide-ranging data, as illustrated by the case in Appendix B.3.
@@
To investigate how code-form plans compare to natural language plans, we conduct comparative experiments by replacing the code plans in our dataset with natural language counterparts [^3] We generate natural language plans using the pipeline in §3.1 with a modified prompt in Appendix A.3..
Figure 4: Comparing natural language planning with CodePlan. The scores of each type of task are averaged across all corresponding benchmarks.
Figure 4 shows the results. Training with natural language plans also leads to moderate improvement over the vanilla baseline. However, CodePlan consistently outperforms its natural language counterparts across all benchmarks by substantial margins. The performance gaps are particularly pronounced on complex reasoning tasks requiring structured planning, such as mathematical reasoning, symbolic reasoning, instruction-following, and decision-making tasks, with relative improvements of 4%, 27.2%, 6.5%, and 27.5%, respectively. This validates the superiority of code representations over natural language.
Table 4: The negative log-likelihood of generating responses in different plan forms. Stage 1/2 refers to planning/surface realization, respectively. “Overall” is calculated by summing up the two parts.
Method Stage 1: $-\log p(Z^{*}|X)$ Stage 2: $-\log p(Y|X,Z^{*})$ Overall
@@
Additionally, we are curious about why planning in code outperforms natural language. We evaluate the performance of the two-stage process: planning and surface realization. For each plan form, Table 4 reports the model’s negative log-likelihood (NLL) on a 10K subset of the training data. The vanilla baseline without explicit planning exhibits a high overall NLL, reflecting its difficulty in directly modeling the complex mapping from prompts to final responses. While natural language planning substantially reduces NLL for surface realization, it introduces a significant challenge in planning compared to CodePlan. We attribute this to two reasons: (1) code provides a more structured and precise representation of complex reasoning logic compared to natural language, thus offering more concise and easier-to-learn plan labels. (2) by framing plan generation as code generation, LLMs can better leverage their pre-training knowledge, as code is far more prevalent in pre-training corpora than natural language plans. Consequently, CodePlan achieves a more substantial overall NLL improvement, providing empirical validation for our theoretical framework in Eq. 4 that requires minimizing the combined difficulty of planning and surface realization.
@@
Prior work primarily focuses on using executable code to directly solve mathematical reasoning tasks (Gao et al., 2023). However, we posit that this approach is inherently limited when tackling broad, multi-step reasoning challenges that necessitate deep natural language understanding capabilities. To substantiate this claim, we conduct comparative experiments by replacing the code plans in our dataset with executable code solutions. Specifically, instead of constructing pseudocode plans, we directly translate each response into executable code that can output the answer after execution with a code interpreter, using the prompt in Appendix A.3. We refer to this baseline as CodeReason.
As presented in Figure 5, while CodeReason achieves the highest performance on SVAMP, a benchmark comprised of relatively simple math word problems, it consistently underperforms CodePlan across a diverse array of reasoning benchmarks that demand extensive natural language understanding capabilities, such as MATH, MuSiQue, and HotpotQA. Notably, CodeReason even lags behind the vanilla training baseline on MATH and MuSiQue. In contrast, CodePlan’s novel framework of generating code plans as an intermediate representation seamlessly integrates robust planning capabilities with the rich language understanding abilities of LLMs, yielding superior performance on intricate multi-step reasoning challenges.
@@ 4 Related Work
Endowing LLMs with robust reasoning abilities remains a formidable challenge. Existing approaches predominantly fall into three categories: (1) Prompting techniques, which use expert-designed prompts to elicit reasoning skills without training (Wei et al., 2022b; Press et al., 2023; Imani et al., 2023; Hong et al., 2024) (2) Task-specific fine-tuning, which curates tailored fine-tuning data or rewards to improve reasoning in specific tasks like mathematical reasoning (Yu et al., 2023; Mitra et al., 2024; Shao et al., 2024; OpenAI, 2024), code reasoning (Le et al., 2022; Shen et al., 2023), instruction-following (Cui et al., 2023), visual reasoning (Cheng et al., 2024a), and decision-making (Zeng et al., 2023; Guan et al., 2024) tasks. However, these approaches often falter in generalizing beyond their intended tasks. (3) Tool integration, which seeks to augment LLMs’ reasoning capabilities by incorporating external tools like calculators (Schick et al., 2024), retrievers (Asai et al., 2024), and code interpreters (Gao et al., 2023). Compared to these works, CodePlan focuses on enhancing LLMs’ high-level planning capabilities in diverse domains, without relying on task-specific prompts, fine-tuning data, rewards, or tools.
@@
Enhancing LLMs’ planning capabilities is crucial for complex reasoning tasks (Yang et al., 2023). Prior work primarily focuses on teaching LLMs to plan in natural language via task-specific prompts (Wang et al., 2023; Khot et al., 2022) or curated fine-tuning data (Yin et al., 2024; Guan et al., 2024). In contrast, CodePlan innovatively introduces code as a structured and versatile plan representation. Recent works also attempt to learn implicit planning (e.g., latent code or verbal words) from wide-ranging text corpora (Zelikman et al., 2024; Cornille et al., 2024). However, these approaches may struggle to automatically unveil effective planning signals from the vast space and often introduce significant computation overhead during training due to online sampling over prior and posterior distributions. In contrast, CodePlan introduces neglectable computation cost as illustrated in Appendix B.2. Additionally, recent works also explore multi-path planning (Yao et al., 2024) and iterative plan refinement (Shinn et al., 2024), which are orthogonal to our work. We leave integrating CodePlan with such techniques for future work.
@@
Recent works have explored leveraging code to empower LLMs for complex reasoning. One approach directly employs prompting techniques (Gao et al., 2023; Ye et al., 2023) or curated fine-tuning data (Gou et al., 2023; Zhou et al., 2023) to generate executable code as a surrogate for natural language response, subsequently utilizing a code interpreter to derive the answer. Despite the precision afforded by executing code, this framework suffers from limited data scalability and is significantly limited to narrow domains such as mathematical calculation. Beyond direct problem-solving, code has also been utilized to enhance LLMs’ capabilities in handling structured reasoning tasks, such as graph generation (Madaan et al., 2022), event structure prediction (Wang et al., 2022), and decision-making (Wang et al., 2024) tasks. Sharing a similar motivation, our work leverages code to represent intricate reasoning structures.
@@ 5 Conclusion
In this work, we introduce a pioneering framework to endow LLMs with robust planning capabilities through the explicit supervision of code-form plans. By reframing plan generation as a code generation task, CodePlan harnesses the structured and versatile nature of code to capture the rich semantics and control flows underpinning sophisticated reasoning processes. Importantly, CodePlan allows the automatic extraction of code-form plans from massive, wide-ranging text corpora without the need for curated, task-specific datasets. We demonstrate the effectiveness of this framework by training CodePlan on a large-scale dataset comprised of 2M natural language problems paired with their corresponding code-form plans. Across an extensive evaluation spanning 13 challenging multi-step reasoning benchmarks, CodePlan demonstrates remarkable efficacy, consistently and substantially outperforming vanilla training by a substantial margin. In-depth analyses further corroborate CodePlan’s increasing performance gain on complex problems and generalization ability. This work paves the way for several promising research directions, including exploring diverse posterior distributions over plans, enabling multi-path planning, facilitating plan verification, reflection and refinement, and realizing agents that can seamlessly leverage external knowledge sources and APIs within their planning and reasoning processes.
@@ A.1 Training Data Curation Details
The curation of high-quality training data is crucial for the success of our approach. To ensure the integrity and relevance of the generated code-form plans, we filter examples without valid terminations in the plans, as this is essential for maintaining logical coherence. Figure 6 illustrates the pipeline for curating the training data.
Figure 6: Training data curation pipeline.
@@ A.2 Data Statistics
Table 6: Number of test examples in the evaluation benchmarks.